OpenStudy (anonymous):

Which of the following lines has a slope of -1/2? x + 2y = 0 x - 2y = 0 -x + 2y = 0

OpenStudy (anonymous):

Rearrange the equations to the general formula of a line, y= mx+c where m is the gradient For x+2y = 0, 2y= -x y= -1/2 x First option is the answer. Im just gonna work out the rest, to let you see how it works For x- 2y = 0 2y = x y= x/2 The gradient in this case is 1/2 For - x+2y = 0 2y = x y= x/2 Same here the gradient is 1/2